### v1

@@ 4.4 Trajectory tracking
In this section, we detail the control system designed to track the trajectories generated in Section 4.3. We use a distributed approach in which each platform has its independent trajectory tracker (see Section 4.4.1). Then, a coordination module, detailed in Section 4.4.2, ensures the synchronization of their movements.
4.4.1 UAV and UGV trajectory tracker
We assume that both the UAV and the UGV accept holonomic velocity commands. We have implemented a simple trajectory tracking scheme based on proportional control of each one of the coordinate axes (see Fig. 11). Our trajectory tracker acts similarly in both UAV and UGV platforms, but adding a proportional height controller in the case of the UAV. Finally, we incorporate an additional yaw control in the UAV, which is used when reaching the inspection point to ensure that the vehicle is pointing towards the area of interest.
@@ 4.4.1 UAV and UGV trajectory tracker
Regarding the magnitude of the velocity command vector ($\boldsymbol{v}$) in the lateral motion, we make use of trapezoidal velocity profiles for waypoint tracking. The maximum velocity for the trapezoid is set to the planned $\rho_{vg}$ and $\rho_{va}$ respectively, but the actual mean velocity is slightly less than the commanded due to the velocity profile.
Finally, the maximum commanded speed has been limited to match the speed of the slowest subsystem, which in this implementation is the UGV. This has been done for safety reasons and to facilitate the coordination of the vehicles. In the experiments, we set the speeds of both the UGV and the UAV ($\rho_{vg}$ and $\rho_{va}$, respectively) to $0.25m/s$.
4.4.2 UGV and UAV coordination
In order to track the trajectory computed by our approach, the UGV and the UAV should follow the trajectories coordinately. This task is carried out by our Trajectory Coordination Module (TCM) (see Fig. 7). In it, we opt for a loosely-coupled solution based on time synchronization. TCM can only command new waypoints to both the UGV and UAV simultaneously, provided that both platforms have successfully reached the previous one. Thus, when a subsystem reaches its current waypoint, it waits for the other in case of need.
5 Experiments
We carried out different field experiments with our marsupial configuration in three different buildings of the UPO, Seville (Spain). Scenario 1 is located in Building 45, home of the Service Robotic Laboratory facilities. Next, we conducted additional tests in two old, abandoned buildings on the same university campus, both of which show significant structural damage. Scenario 2 is situated in the Abandoned Thermal Station, while Scenario 3 is located in the Old Theater building. Additional details of the experiments are provided in the video attached to this paper.
5.1 Scenario 1: Flight Duration Tests
The main goal of this experiment was to perform the necessary system integrity checks of the whole marsupial system and to estimate the flight endurance of the system in manual mode. To obtain a valid estimation, we mimicked the stop-and-go behavior in an inspection mission, obtaining a total trajectory length exceeding one kilometer on both the UAV and the UGV. Figure 12 shows two snapshots obtained during the experiment.
@@ 5.1 Scenario 1: Flight Duration Tests
Figure 12: Two snapshots of the flight duration tests at building 45.
The experiment exceeded one hour duration, which allowed us to get information about the discharging rate of both the onboard UAV and the batteries on the UGV used for powering the UAV via the Elistair tether system. Table 1 represents the level of the batteries powering the tether system and the onboard UAV battery used as a backup, for ease of comparison, during the experiment. In the second row, we can find the battery percentage of the onboard battery used for backup purposes on the UAV. Note that this battery slowly discharges until stabilizing at 86%, when the voltage level of the battery reaches the one provided by the Elistair system. Regarding the batteries onboard the UGV, their discharge rate remains constant during the experiment and very close together. At the end of the experiment, the remaining battery level was $53\%$, giving an average power consumption on the UAV of 1635 W.
Table 1: Charge of the batteries, in [%] of the different systems as the mission time progresses.
Battery Percentage
Time [min] 0 15 30 45 60
Onboard UAV Battery [%] 100 92 88 86 86
Battery Pack 1 [%] 100 88 76 64 52
Battery Pack 2 [%] 100 89 77 66 54
5.2 Scenario 2: UAV Autonomous Navigation Test
@@
In Scenario 2, we visited the Abandoned Thermal Station building of the UPO. This building was built in the decade of the 1950s and presented great structural damage and debris on its surface that prevented our UGV ARCO platform from safely performing autonomous navigation on it. Therefore, we designed some experiments focused on the UAV system to test the indoor UAV localization and navigation systems.
5.2.1 UAV Localization tests
@@
In a first set of experiments, we tested the ability of our localization system to provide us with an online estimation of the pose of the UAV in manual mode. To this end, we designed a challenging experiment in which we made our UAV move around the whole experimental scenario at different height levels while recording the data from the sensors. Then, we tested different localization methods to compare their results and their average computation time. To evaluate the precision of the localization algorithm, we got the distance from each LiDAR detection inside a measurement to the closest obstacle in the map, to check for errors in the localization. The lower the average error in a measurement is, the better the vehicle is localized in the environment. This was necessary due to the lack of ground truth from external systems available. Table 2 represents the mean distance from the LiDAR points to the closest obstacle with different localization methods and their average execution time. Results indicate that our DLL method used in our system outperforms the MCL in terms of precision and runs significantly faster when compared to the ICP method, achieving similar errors. Moreover, the measured distance from the closest obstacle results indicate that the localization error should be in the order of centimeters, while also inferring a good precision in the estimation of the yaw. This precision is enough to safely perform cooperative missions in our marsupial system. Thus, DLL is a very convenient alternative for having precise localization when performing long-endurance inspection missions.
@@
Table 2: Average computation time of the update step on each method and mean distance from each point of the point-cloud (PC) to the closest obstacle in the map in Scenario 2.
Method Avg. Computation time (s) Avg. PC distance to obstacle (m)
AMCL-3D 0.54 0.45
ICP 2.67 0.13
DLL 0.08 0.14
5.2.2 UAV autonomous flight tests
@@
Once we demonstrated that our localization system was able to obtain a precise estimation of the pose of the robot in real-time, we executed several autonomous tests for the UAV to test our trajectory tracking system. To this end, a simple inspection plan was designed and generated the trajectory to be tracked by the UAV by using our trajectory planner, assuming a static UGV. Figure 13 represents the generated trajectories to perform the UAV autonomous navigation test represented in the 3D model of the environment. Our system was able to autonomously follow the trajectory, obtaining mean lateral cross-track and height errors of 0.07 m and 0.08 m, respectively.
@@
Figure 13: Reference trajectory (blue line) tracking experiments in Scenario 2. The position of the reel system is on a green dot and the red lines represent the tether configurations during the experiment.
5.3 Scenario 3: Emulated Inspection Test
@@
Finally, to test the whole system in autonomous mode, we emulated a structural assessment inspection. In Scenario 3, our marsupial system is configured with a mission designed to look for emulated defects in the Old Theater building of the UPO. To emulate the defects, we disposed of an array of twelve AR markers with a side of fifteen centimeters on one of the walls of the building (see Fig. 1b). We equipped the UAV with a high-resolution camera to automatically detect the markers by using the ArUco library [43].
@@
Figure 8.a represents an obtained 3D map of the environment, which was used for localization and path planning purposes on both platforms in Scenario 3. The goal of the experiment is to make an in-depth inspection of one of the walls of the building. To this end, we disposed of the PoIs of our inspection plan following a zig-zag line. Then, we generated a trajectory for the marsupial system using our trajectory planner module (see Section 4.3).
In the experimental session, the batteries of the platform allowed us to perform the execution three times without the need of changing any batteries, for a total flight time of about fifty minutes. The trajectories commanded to the UAV and the UGV are represented in Fig. 14. The UAV platform accurately tracked the trajectories, detecting all emulated defects in three experimental runs, as illustrated in Fig. 15. Statistics on the error in the estimated position of the detected ArUco markers are provided in Table 3. In particular, the mean detection error of each marker is on the order of centimeters. Moreover, the cases with greater errors are usually related to a low number of observations of the AR marker.
@@
Figure 14: Computed trajectory for the inspection mission of UGV (in green dots) and UAV (in blue dots). The square and the sphere in white color represent the initial and the final positions, respectively.
Figure 15: Current and past detections of emulated defects obtained in the inspection mission of Scenario 3.
@@ 6 Discussion and take-home messages
This paper presents the design, implementation, and validation of a marsupial robotic system capable of performing long-duration inspections in GNSS-denied environments. Experimental results confirm the hypothesis that a tethered UAV-UGV system can significantly extend UAV flight time, opening the door to long-duration inspections. The accuracy in detecting AR markers, with a mean error in the order of centimeters, demonstrates the viability of the system for detailed inspection tasks.
@@
Even though power consumption was higher than expected, the platform proved capable of performing inspection missions lasting more than two hours. The additional discharge may be due to extra connected components which were not included in the calculations, such as the LiDAR and the router, or due to heat losses in the different voltage conversions. However, the flight duration can be further extended by adding more battery packs to the UGV, as we are far away from its payload limits.
A summary of the most valuable lessons we have learned from experimentation is mentioned in the Section 6.1. Then, Section 6.2 lists the advantages and disadvantages of our approach compared to other alternatives.
6.1 Lessons Learned
Several factors are essential for the success of a multi-robot experiment. In this section, we highlight some key insights that we have acquired from our experimentation.
  • Design a setup procedure and a checklist to ensure experimental quality. When performing experiments with complex systems such as our tether marsupial system, it is important to carefully design a procedure for initializing and checking the systems before starting a mission. We follow the steps outlined in Algorithm 1. In Steps 1–3, the UGV is started system first, as its onboard computer hosts the ROS master node required for system-wide communication. Additionally, we power up the UAV as late as possible (Step 8) to maximize its available flight time. Conducting integrity tests on each onboard sensor is essential to avoid incomplete experiment recordings. Specifically, we verify that each sensor’s data rate and quality meet expected standards—this is done in Step 4 for the UGV and Step 10 for the UAV.
  • Avoid dangerous configurations of the UAV, the UGV and the tether. The Elistair LIGH-T V4 station has a DC motor that performs a constant tension to the tether, keeping it taut. Experimentally, we found out that a tension of one Newton is enough to keep it straight in most situations. The main challenge here is that this tension introduces disturbances that must be managed by the UAV’s control system. When tension is applied downwards, it has a similar effect to increasing the system’s payload. However, when tension is applied laterally, the control system must adjust the UAV’s attitude to compensate for the disturbances. Therefore, it is advisable to penalize configurations with low elevation angles. Lastly, we secure the tether as close to the center of gravity of the UAV as possible to prevent the generation of torques.
  • Prepare and label several UAV batteries the day before an experimental session. UAVs need a significant amount of energy to complete long-duration flights, which typically requires fresh batteries for each experimental run. However, LiPO batteries should not be stored fully charged for extended periods. Therefore, it is advisable to check and charge the UAV batteries the day before an experimental session. It also applies to our system because even if the energy consumption of the onboard UAV backup battery is low during the experiment (see Table 1), the M210 firmware expects the onboard batteries to be fully charged to power up the system.
Algorithm 1 Experiment Setup and Execution Checklist
1:Turn on the battery packs and the inverter at the UGV.
2:Turn on router onboard the UGV.
3:Turn on the UGV and start its ROS nodes, including the ROS master process necessary for communications.
4:Turn on laptop at the base station and check for ROS connectivity with the UGV.
5:Perform integrity tests of the sensors and software systems on the UGV.
6:Set the UAV and the UGV at their initial poses.
7:Turn on the Elistair Power Tether.
8:Just before the experiment: Turn on the UAV computer.
9:Check for connectivity to the UAV system
10:Perform integrity tests to the sensors and software systems onboard the UAV.
11:Start the TCM module on the base station.
12:Start the recording of the experiments including video and data logging.
13:Start the mission.
6.2 Marsupial systems: benefits and drawbacks
With all the knowledge gathered from experimentation, we are confident that our marsupial system is a viable alternative that has multiple advantages over existing approaches in the literature. Thus, the following list summarizes the benefits of our approach.
  • The UGV can supply power to the UAV via the cable, significantly extending flight time compared to the UAV’s limited battery capacity. UAV time fly can extend over one hour.
  • Unlike UAVs tethered to a fixed location, the UGV provides mobility, enabling the UAV to inspect a wider range of areas of interest.
  • The UAV offers aerial mobility and monitoring capabilities, while the UGV provides a stable power source and additional payload capacity.
  • The cable can serve as a reliable communication link and can also enhance safety by providing a physical connection between the two robots.
  • Marsupial UAV-UGV systems have proven valuable in hazardous and confined environments, as exemplified by their use by the CSIRO team during the DARPA Subterranean Challenge. Our system has been able to carry out a long-endurance inspection task, in a GNSS-denied environment, with accurate localization and safe trajectory planning.
However, even though our approach has been demonstrated as an effective alternative to perform long-duration inspection tasks, there are some disadvantages compared to other existing approaches. Besides, our approach has room for improvement in some areas.
  • The tether can introduce significant disturbances in the UAV control subsystem. Interactions between the tether and the UAV generate dynamic forces that affect stability and control.
  • The tether restricts the UAV’s range of motion and requires careful management to prevent collisions. Maintaining proper tension is crucial: downward tension impacts the system payload, while lateral tension requires UAV control adjustments to compensate for perturbations.
  • There is a risk of the tether becoming entangled or snagged on obstacles.Therefore, the mission should be carefully planned to prevent the cable from touching any obstacles. In addition, the cable should also monitored during the experiments, which should be canceled in the event of entanglements or other unexpected events.
  • Variable speed trajectory tracking can be implemented to smoothen the coordinated movement of the UAV and the UGV.

### v2

@@ 4.4 Trajectory Tracking
In this section, we detail the control system designed to track the trajectories generated in Section 4.3. We use a distributed approach in which each platform has its independent trajectory tracker. Then, a coordination module ensures the synchronization of their movements.
Thus, for trajectory tracking, we assume that both the UAV and the UGV accept holonomic velocity commands. We have implemented a simple trajectory tracking scheme based on proportional control of each one of the coordinate axes (see Figure 11). Our trajectory tracker acts similarly in both UAV and UGV platforms but adds a proportional height controller in the case of the UAV. Finally, we incorporate an additional yaw control in the UAV, which is used when reaching the inspection point to ensure that the vehicle is pointing towards the area of interest.
@@
Regarding the magnitude of the velocity command vector ($\boldsymbol{v}$) in the lateral motion, we make use of trapezoidal velocity profiles for waypoint tracking. The maximum velocity for the trapezoid is set to the planned $\rho_{vg}$ and $\rho_{va}$, respectively, but the actual mean velocity is slightly less than the commanded due to the velocity profile.
The maximum commanded speed has been limited to match the speed of the slowest subsystem, which in this implementation is the UGV. This decision was made for safety reasons and to facilitate the coordination of the vehicles. In the experiments, we set the speeds of both the UGV and the UAV ($\rho_{vg}$ and $\rho_{va}$, respectively,) to 0.25 m/s.
Finally, in order to track the trajectory computed by our approach, the UGV and the UAV should follow the trajectories coordinately. This task is carried out by our TCM (see Figure 7). We opt for a loosely coupled solution based on time synchronization. TCM can only command new waypoints to both the UGV and UAV simultaneously, provided that both platforms have successfully reached the previous one. Thus, when a subsystem reaches its current waypoint, it waits for the other in case of need.
5 Experimental Results
We carried out different field experiments with our marsupial configuration in three different buildings of the UPO, Seville (Spain). Scenario 1 is situated in Building 45, home to the Service Robotics Laboratory facilities. Next, we conducted additional tests in two old, abandoned buildings on the same university campus, both of which show significant structural damage. Scenario 2 takes place within the Abandoned Thermal Station, whereas Scenario 3 is set in the Old Theater building. The logs of all experiments are accessible through our institutional repository (https://robotics.upo.es/datasets/marsupial, accessed on 3 Nov 2025). Furthermore, additional experimental details are presented in the video accompanying this paper.
5.1 Scenario 1. Experiments 1–3: Flight Duration Tests
The main goals of this experiment were to perform the necessary system integrity checks of the whole marsupial system and to estimate the flight endurance of the system. We performed three different experiments (Experiments 1–3) in which our marsupial system was commanded in manual mode, mimicking the stop-and-go behavior in an inspection mission. The duration of all the experiments exceeded one hour, with Experiment 3 exceeding two hours. All the experiments were performed in the same load condition. Figure 12 shows two snapshots obtained during the experiments.
@@ 5 Experimental Results
Figure 12: Two snapshots of one flight duration test at building 45.
Figure 13 represents the level of the batteries powering the LTS4 and the onboard UAV backup battery in Experiments 1–3. Remarkably, the discharge rate of the LTS4 battery remained approximately constant throughout all three experiments; the discharge is 46% in one hour in the worst case. Regarding the backup battery, the remaining level was above 80% in all three cases, which indicates that there was not significant usage, as the LTS4 system worked without issues during all three experiments. Moreover, this remaining battery level would allow the monitoring operator to safely land the UAV in the event of an LTS4 malfunction, as was designed in Section 2.4. With all the gathered data, we can conclude that our proposed system can perform inspection missions lasting up to two hours.
Figure 13: Remaining battery levels of LTS4 and backup batteries in Experiments 1–3.
5.2 Scenario 2: Abandoned Thermal Station
@@
Scenario 2 is located at the Abandoned Thermal Station building of the UPO, built in the decade of the 1950s and currently exhibiting significant structural damage and surface debris. These conditions prevented our ARCO platform from autonomously navigating within the building. Consequently, we designed a set of experiments focused on the UAV system to evaluate its indoor localization and navigation capabilities.
5.2.1 Experiments 4–5: UAV Localization Tests
@@
In Experiments 4 and 5, we tested the precision of our localization system in manual mode. To this end, we designed a challenging experiment in which we made our UAV move around the whole experimental scenario at different height levels while recording the data from the sensors. Then, we tested different localization methods to compare their results and their average computation time. To evaluate the precision of the localization algorithm, we obtained the distance from each LiDAR detection inside a measurement to the closest obstacle in the map to check for errors in the localization. The lower the average error in a measurement is, the better the vehicle is localized in the environment. This was necessary due to the lack of ground truth from external systems available. Table 3 represents the mean distance from the LiDAR points to the closest obstacle with different localization methods and their average execution time. We configured the MCL method with 500 particles, and we considered that each measure of the OS1 sensor has a standard deviation of 0.1 m and that the odometry has a relative error of 10%. We used the ICP implementation of the Point Cloud Library (PCL) with a maximum of 50 iterations, a maximum correspondence distance of 0.1 m and a RANSAC outlier rejection threshold of one meter. Finally, our DLL solver was implemented with the Ceres library, being configured to use 8 threads. All the experiments were executed in the NUC11TNKi5 computer onboard the UAV, which has a 13th Gen Intel® Core™ processor with 12 cores, one thread per core, and 16 GB of RAM.
Results indicate that our DLL method outperforms the MCL in terms of precision and runs significantly faster when compared to the ICP method, achieving similar errors. While these results offer an indirect evaluation of the accuracy of the localization, in practice we found that it was enough to safely perform cooperative missions in our marsupial system, making DLL the better method for real-time localization.
For more details about the experiment, please refer to the attached video, where you can check the behavior of our DLL in the localization tests.
@@
Table 3: Average computation time of the update step on each method and average distance from each point of the point-cloud to the closest obstacle in the map in Scenario 2.
Exp. Method Comp. time (s) Avg. Dist. (m)
MCL-3D 0.71 0.45
4 ICP 1.8 0.13
DLL 0.07 0.14
MCL-3D 0.64 1.25
5 ICP 2.9 0.56
DLL 0.11 0.26
5.2.2 Experiment 6: UAV Autonomous Flight Test
@@
After demonstrating that our localization system could accurately estimate the robot’s pose in real time, we conducted Experiment 6 to evaluate the UAV’s trajectory tracking system. To this end, a simple inspection plan was designed, and the trajectory to be tracked by the UAV was generated using our trajectory planner, assuming a static UGV. Figure 14 represents the generated trajectory to perform the UAV autonomous navigation test represented in the 3D model of the environment. The tracking system was able to autonomously follow the trajectory in all the executions, achieving mean lateral cross-track and altitude errors of 0.07 m and 0.08 m, respectively.
Please refer to the attached video for additional details regarding the performance of the trajectory tracker.
@@
Figure 14: Reference trajectory (blue line) tracking experiments in Scenario 2. The position of the reel system is on a green dot, and the red lines represent the tether configurations during the autonomous flight test. The points in white represent the 3D map of Scenario 2.
5.3 Scenario 3. Experiment 7: Emulated Inspection Test
@@
Finally, to test the whole system in autonomous mode, we emulated a structural assessment inspection in Experiment 7, which was carried out in Scenario 3. In Experiment 7, our marsupial system is commanded to execute a mission designed to look for emulated defects in the Old Theater building of the UPO. To emulate the defects, we deployed an array of twelve Augmented Reality (AR) markers with a side of fifteen centimeters on one of the walls of the building (see Figure 1b). We equipped the UAV with a high-resolution camera to automatically detect the markers by using the ArUco library [43].
@@
Figure 8a represents an obtained 3D map of the environment, which was used for localization and path planning purposes on both platforms in Scenario 3. The goal of the experiment is to make an in-depth inspection of one of the walls of the building. To this end, we distributed the PoIs of our inspection plan following a zig-zag line. The distribution of the PoIs was sufficient to ensure complete wall coverage with a spatial resolution of twenty pixels per square centimeter, which is adequate for the detection of the emulated defects. We used the Fields2Cover library [44] to generate the sequence of PoIs (see Figure 15).
@@
Figure 15: Computed trajectory for the inspection mission of UGV (in green dots) and UAV (in blue dots). The square in white represents the initial UAV position. The spheres in white represent the PoIs to be visited.
5.3.1 Trajectory Planning Results
With the inspection plan defined as a sequence of PoIs, we generated a trajectory for the marsupial system using our trajectory planner module (see Section 4.3). Note that we need multiple calls to the planner to generate trajectories connecting the different PoIs.
Table 4 presents the statistics of our trajectory planning method compared to the approach proposed in [28], which accounts for a non-straight tether when generating the complete inspection mission. Both methods successfully generated trajectories. Notably, our planning method achieves lower execution times compared to the baseline.
Regarding clearance, our method yields a smaller mean distance between the tether and both the obstacles and the UAV. This is expected, as we consider only a straight tether, which reduces the available maneuvering space. Finally, in terms of total traveled distance, both methods produce similar UAV paths; however, our method requires the UGV to move slightly more due to the LoS constraint.
In conclusion, our method is capable of generating safe trajectories for the system with lower computation time compared to the approach in [28].
Table 4: Trajectory planning statistics for the proposed planner (LoS) and the planner proposed in [28] (catenary).
Method Execution time (s) Avg. distance to obstacles (m) Trajectory Length (m)
Cable UAV UGV UAV UGV
Catenary 3.37 1.51 1.54 1.09 50.40 11.59
LoS 0.49 1.11 1.08 1.65 48.30 14.14
5.3.2 Mission Execution Results
In the experimental session, our battery-powered LTS4 system allowed us to execute the inspection mission three times without the need of changing any batteries, for a total duration of thirty-five minutes. The trajectories commanded to the UAV and the UGV are represented in Figure 15 and are composed of 245 waypoints. As detailed in Section 2.4, a safety operator supervised the execution of the experiment, ensuring the absence of unexpected tether entanglements and monitoring the charge level of the UAV’s onboard backup battery. Figure 16 illustrates the battery levels of the UAV, the UGV, and the batteries supplying the LTS4 system. The backup battery level remained consistently above 80% throughout the experiment. Moreover, since no unexpected entanglements were observed, no operator intervention was required. Regarding the UGV’s onboard batteries, their voltage level remained approximately constant at 25.5 V during the entire experiment. Although voltage is not a precise indicator of the remaining charge for LiFePO4 batteries, this stability suggests that the battery operated within its mid-range (approximately 30 to 70%). Finally, the batteries powering the LTS4 maintained a charge level above 60%, indicating that safe operation could have continued for more than one hour after the experiment.
The performance of our trajectory tracking systems was also remarkable. The system was capable of reaching all the planned waypoints in all three experimental runs. Finally, we evaluate the performance of the TCM. Table 5 presents the statistics of the waiting times for both the UAV and the UGV during each mission execution. The results are highly consistent. The UAV experiences a higher number of waiting events due to its lower velocity. However, the difference in waiting times is small, averaging approximately 1.5 s. Although this value may seem relatively high, it should be noted that the trajectory includes segments where only one vehicle is moving, causing the other to wait until the next waypoint is reached. Finally, the relatively high standard deviation indicates the presence of outliers in which the waiting times are prolonged for several seconds. This typically occurs at the beginning of the experiment.
Figure 16: Remaining battery levels of the LTS4 (red and blue) and backup batteries and measured voltage of the UGV batteries (green and cyan) in Experiment 7.
Table 5: Average waiting time and standard deviation for the UAV and UGV induced by the TCM on Experiment 7. Values in parentheses indicate the number of waiting events.
Mission execution UAV waiting time (s) UGV waiting time (s)
1 1.89 $\pm$ 2.21 (147) 1.51 $\pm$ 1.93 (147)
2 1.43 $\pm$ 2.73 (144) 1.43 $\pm$ 2.14 (101)
3 1.47 $\pm$ 2.39 (142) 1.39 $\pm$ 1.78 (103)
The marsupial system was able to detect all the emulated defects in the three experimental runs, as illustrated in Figure 17. Statistics on the error in the estimated position of the detected ArUco markers are provided in Table 6. In particular, the mean detection error of each marker is on the order of centimeters. Moreover, the cases with greater errors are usually related to a low number of observations of the AR marker. The achieved precision of the estimated position of the AR markers also validates the precision of our localization system.
Figure 17: Current and past detections of emulated defects (AR Markers) obtained in Scenario 3, Experiment 7. (a) Current image from the onboard camera with an overlay of the current detection (ID 413). (b) Estimated positions of the UAV, the UGV and the current and past detections over the 3D map of the environment.
@@ 6 Discussion
This paper has presented the design, implementation, and validation of a marsupial robotic system capable of performing long-duration inspections in GNSS-denied environments. Experimental results confirm the hypothesis that a tethered UAV-UGV system can extend UAV flight time to more than two hours of continuous operation, opening the door to long-duration inspections. The accuracy in detecting AR markers, with a mean error in the order of centimeters, demonstrates the viability of the system for detailed inspection tasks.
@@
The platform proved capable of performing inspection missions lasting more than two hours, which aligns with our estimates when accounting for the power consumption of the UGV’s internal LiDAR and router, both of which are connected to the inverter. However, the flight duration could be further extended by integrating more battery packs to the UGV, as the platform remains well below its payload limit.
With the knowledge gathered from our experimentation, we are confident that the proposed marsupial system represents a viable alternative with several advantages over existing approaches in the literature. One of its key strengths lies in the capability of the UGV to supply power to the UAV through the tether, which significantly extends the UAV’s endurance beyond the limitations imposed by its onboard battery. In practice, this enables flight times of over two hours, a remarkable improvement over conventional untethered UAVs. Furthermore, in contrast to systems where the UAV is tethered to a fixed station, our mobile UGV base provides additional flexibility by enabling the UAV to cover a broader range of inspection areas. The complementary nature of both platforms enhances mission capabilities: while the UAV contributes aerial mobility and monitoring functionality, the UGV acts as a stable power source and offers additional payload capacity. Marsupial UAV–UGV systems have already demonstrated their utility in hazardous and confined environments, as shown, for instance, by the CSIRO team in the DARPA Subterranean Challenge [19] and the Oxpecter platform [21]. Our approach significantly advances autonomy, demonstrating long-endurance inspection in GNSS-denied environments with precise localization and safe trajectory planning.
Despite these benefits, our approach is not without limitations. While it is true that tethering limits the UAV’s maneuverability and may constrain its flyable space, requiring careful management to prevent collisions, the extended flight duration enables our solution to carry out long-duration inspection missions, where steady flight is essential to ensure the validity of the measurements. In our view, the extended flight time outweighs the reduction in maneuverability and speed for the applications considered. Maintaining proper tension is essential, since downward tension directly affects payload capacity, whereas lateral forces demand corrective control actions from the UAV to maintain stability. However, we found that by configuring LTS4 to perform a tension of 1 N to the UAV, we can keep the cable straight when only increasing the required force to hover in a 2. 5%, as the UAV weighs approximately four kilograms.
Another drawback is that the tether carries the risk of becoming entangled or snagged on environmental obstacles. To mitigate this, missions are carefully designed to prevent any contact between the tether and the environment. The cable is continuously monitored during the experiments, which are halted if entanglements or other unexpected events occur. Finally, while our system achieves effective cooperation between the UAV and UGV, further improvements could be obtained through variable-speed trajectory tracking, which would enable smoother coordination between both platforms.